# Smoothed Analysis of Online Metric Matching with a Single Sample: Beyond Metric Distortion

### v3

@@ 3 Random-Subtree Algorithm
The Random-Subtree (RS) algorithm proposed by [DBLP:conf/icalp/GuptaL12] is a natural starting point for online matching on HST metrics. In the adversarial setting, it achieves an $O(\log\Delta)$ competitive ratio for $\Delta$-ary $\alpha$-HST metrics, provided $\alpha=\Omega(\log\Delta)$. Here we revisit the RS algorithm in the stochastic setting where servers and requests are independently drawn from a distribution $\mathbb{D}=\prod_{i=1}^{n}\mathbb{D}_{i}$. Later on, we will translate the guarantees to the setting where servers are adversarial and requests are drawn from $\mathbb{D}$ by applying Lemma 2.5. Rather than its competitive ratio, we upper bound the cost of the $\mathrm{RS}$ algorithm for arbitrary $\alpha\geq 2$.
@@
The proof of Lemma 3.2 largely follows the proof strategy of [DBLP:conf/icalp/GuptaL12, Theorem 4.1], with two key differences. First, their analysis assumes $\alpha=\Omega(H_{\Delta})$, ensuring the HST is sufficiently well-separated so that the costs incurred at lower levels are dominated by those at higher levels. In contrast, we do not rely on this separation, and hence the costs incurred at lower levels must be handled explicitly. Second, our proof is relatively simpler: since we only bound the absolute cost of the algorithm, we avoid the more involved step of characterizing the offline optimum, which is required in [DBLP:conf/icalp/GuptaL12].
@@
Lemma 3.3 (Lemma 4.3 in [DBLP:conf/icalp/GuptaL12]).
@@
We recall the following useful decomposition of $\mathrm{cost}(\mathrm{RS};S,R\cup R^{\prime})$ from [DBLP:conf/icalp/GuptaL12].
Lemma 3.4 (Fact 4.8 in [DBLP:conf/icalp/GuptaL12]).
@@ 4 BBGN Algorithm
In this section, we turn from the $\mathrm{RS}$ algorithm to the algorithm proposed by [DBLP:journals/algorithmica/BansalBGN14], henceforth referred to as the BBGN algorithm, which outperforms the $\mathrm{RS}$ algorithm in certain parameter regimes. The BBGN algorithm is $O(\log n)$-competitive for $\alpha$-HST metrics with any constant $\alpha>1$ when both $S$ and $R$ are adversarial. Our interest, however, lies in analyzing its cost when both $S$ and $R$ are drawn from $\mathbb{D}$. As before, we will later translate these guarantees to the setting where servers are adversarial and requests are drawn from $\mathbb{D}$ by applying Lemma 2.5.
@@
The rest of this section is devoted to proving Theorem 4.1. For each leaf node $v$ of the HST and $\ell\in\{0,1,\ldots,h\}$, let $p(v,\ell)$ be the ancestor of $v$ with height $\ell$. The $\mathrm{BBGN}$ algorithm was originally developed in a restricted reassignment online model, which permits limited reassignment of previously arrived requests, and then adapted to the standard online model. We recall here a bound for the cost of the $\mathrm{BBGN}$ algorithm that applies to all $S$ and $R$, obtained by combining [DBLP:journals/algorithmica/BansalBGN14, Lemmas 4.2, 4.3, and 4.4], and we refer to [DBLP:journals/algorithmica/BansalBGN14] for a more formal description of the algorithm.
Lemma 4.2 ([DBLP:journals/algorithmica/BansalBGN14]).
@@
It is well known that min-cost perfect matchings in HSTs admit a useful characterization (see, e.g., [DBLP:journals/algorithmica/BansalBGN14, Lemma 4.1]): every min-cost perfect matching between $S$ and $R$ can be obtained by a greedy algorithm that repeatedly matches the closest request-server pair and then recurses on the remaining instance. Equivalently, the number of matched pairs whose two endpoints lie in different subtrees of a given internal node can be expressed as follows.
@@ 5 Euclidean Metrics
Lemma 5.7 (Lemma 14 in the arXiv version of [yang2025online]).
@@ 6 Discussion and Future Directions
In this paper, we study the online metric matching problem for the Euclidean space $0,1]^{d}$ when servers are adversarial and requests are independently drawn from distinct smooth distributions. We present an $O(1)$-competitive algorithm for $0,1]^{d}$ with $d\neq 2$, given a single sample from each request distribution. A key feature of our approach is that, by directly upper-bounding the algorithm’s cost after a simple deterministic metric embedding, we bypass the $\Omega(\log n)$ competitive-ratio barrier that arises in the adversarial setting due to metric distortion. Since metric embeddings into HSTs have already been proven extremely effective for related online problems such as $k$-server [DBLP:journals/jacm/BansalBMN15, DBLP:conf/stoc/BubeckCLLM18], $k$-taxi [DBLP:conf/soda/GuptaKP24], and several variants of online metric matching [DBLP:conf/stoc/EmekKW16, DBLP:conf/soda/BhoreFT24], a natural and exciting future direction is to extend our techniques to these problems.
Our guarantees rely on requests being independently sampled. An intriguing direction would be to see what forms of correlation among requests might still permit an $o(\log n)$ competitive ratio. As a starting point, recent breakthroughs in smoothed analysis of online learning [DBLP:conf/colt/BlockDGR22, DBLP:journals/jacm/HaghtalabRS24] allow each arrival’s distribution—while required to be smooth—to depend on the realized history of arrivals and algorithmic decisions, and their techniques may extend to our setting. In addition, the correlation models studied in online stochastic matching [DBLP:conf/sigecom/AouadM23] and prophet inequalities [DBLP:journals/teco/ImmorlicaSW23] present additional promising avenues.

### v4

@@ 3 Random-Subtree Algorithm
The Random-Subtree (RS) algorithm proposed by [GL12] is a natural starting point for online matching on HST metrics. In the adversarial setting, it achieves an $O(\log\Delta)$ competitive ratio for $\Delta$-ary $\alpha$-HST metrics, provided $\alpha=\Omega(\log\Delta)$. Here we revisit the RS algorithm in the stochastic setting where servers and requests are independently drawn from a distribution $\mathbb{D}=\prod_{i=1}^{n}\mathbb{D}_{i}$. Later on, we will translate the guarantees to the setting where servers are adversarial and requests are drawn from $\mathbb{D}$ by applying Lemma 2.5. Rather than its competitive ratio, we upper bound the cost of the $\mathrm{RS}$ algorithm for arbitrary $\alpha\geq 2$.
@@
The proof of Lemma 3.2 largely follows the proof strategy of [GL12, Theorem 4.1], with two key differences. First, their analysis assumes $\alpha=\Omega(H_{\Delta})$, ensuring the HST is sufficiently well-separated so that the costs incurred at lower levels are dominated by those at higher levels. In contrast, we do not rely on this separation, and hence the costs incurred at lower levels must be handled explicitly. Second, our proof is relatively simpler: since we only bound the absolute cost of the algorithm, we avoid the more involved step of characterizing the offline optimum, which is required in [GL12].
@@
Lemma 3.3 (Lemma 4.3 in [GL12]).
@@
We recall the following useful decomposition of $\mathrm{cost}(\mathrm{RS};S,R\cup R^{\prime})$ from [GL12].
Lemma 3.4 (Fact 4.8 in [GL12]).
@@ 4 BBGN Algorithm
In this section, we turn from the $\mathrm{RS}$ algorithm to the algorithm proposed by [BBG+14], henceforth referred to as the BBGN algorithm, which outperforms the $\mathrm{RS}$ algorithm in certain parameter regimes. The BBGN algorithm is $O(\log n)$-competitive for $\alpha$-HST metrics with any constant $\alpha>1$ when both $S$ and $R$ are adversarial. Our interest, however, lies in analyzing its cost when both $S$ and $R$ are drawn from $\mathbb{D}$. As before, we will later translate these guarantees to the setting where servers are adversarial and requests are drawn from $\mathbb{D}$ by applying Lemma 2.5.
@@
The rest of this section is devoted to proving Theorem 4.1. For each leaf node $v$ of the HST and $\ell\in\{0,1,\ldots,h\}$, let $p(v,\ell)$ be the ancestor of $v$ with height $\ell$. The $\mathrm{BBGN}$ algorithm was originally developed in a restricted reassignment online model, which permits limited reassignment of previously arrived requests, and then adapted to the standard online model. We recall here a bound for the cost of the $\mathrm{BBGN}$ algorithm that applies to all $S$ and $R$, obtained by combining [BBG+14, Lemmas 4.2, 4.3, and 4.4], and we refer to [BBG+14] for a more formal description of the algorithm.
Lemma 4.2 ([BBG+14]).
@@
It is well known that min-cost perfect matchings in HSTs admit a useful characterization (see, e.g., [BBG+14, Lemma 4.1]): every min-cost perfect matching between $S$ and $R$ can be obtained by a greedy algorithm that repeatedly matches the closest request-server pair and then recurses on the remaining instance. Equivalently, the number of matched pairs whose two endpoints lie in different subtrees of a given internal node can be expressed as follows.
@@ 5 Euclidean Metrics
Lemma 5.7 (Lemma 14 in the arXiv version of [YY25]).
@@ 6 Discussion and Future Directions
In this paper, we study the online metric matching problem for the Euclidean space $0,1]^{d}$ when servers are adversarial and requests are independently drawn from distinct smooth distributions. We present an $O(1)$-competitive algorithm for $0,1]^{d}$ with $d\neq 2$, given a single sample from each request distribution. A key feature of our approach is that, by directly upper-bounding the algorithm’s cost after a simple deterministic metric embedding, we bypass the $\Omega(\log n)$ competitive-ratio barrier that arises in the adversarial setting due to metric distortion. Since metric embeddings into HSTs have already been proven extremely effective for related online problems such as $k$-server [BBM+15, BCL+18], $k$-taxi [GKP24], and several variants of online metric matching [EKW16, BFT24], a natural and exciting future direction is to extend our techniques to these problems.
Our guarantees rely on requests being independently sampled. An intriguing direction would be to see what forms of correlation among requests might still permit an $o(\log n)$ competitive ratio. As a starting point, recent breakthroughs in smoothed analysis of online learning [BDG+22, HRS24] allow each arrival’s distribution—while required to be smooth—to depend on the realized history of arrivals and algorithmic decisions, and their techniques may extend to our setting. In addition, the correlation models studied in online stochastic matching [AM23] and prophet inequalities [ISW23] present additional promising avenues.